# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ 5 Experiments and Results
We present our overall test results in Table 4. Based on the experimental setup presented in Valderrama et al. (2022), we compare the performance of the best TAPIS model with a CNN-based baseline. For this alternative baseline, we employ SlowFast (Feichtenhofer et al. (2019)) as the central video feature extractor and use ISINet (González et al. (2020)) (based on Mask R-CNN (He et al. (2017))) as the region proposal network for short-term tasks. We use the exact implementation details of SlowFast presented in Valderrama et al. (2022). We call this model SlowFast in Table 4. We also compare TAPIS with an alternative implementation using Video Swin Transformer (VST) (Liu et al. (2022)) as the video feature extractor and maintaining our proposed region classification head; we call this model TAPIS-VST in Table 4. The results in Table 4 demonstrate TAPIS’s and VST-TAPIS’s superiority over the CNN-based counterpart. This behavior proves that the inherent sequence-processing capability of transformers allows for a superior comprehension of temporal progressions, and their attention mechanisms result in better recognition of instruments, anatomical cues, and other relevant visual contexts in surgical videos, leading to higher performance in all the proposed semantic tasks. Likewise, TAPIS surpasses TAPIS-VST in most tasks and achieves comparable instrument segmentation performance with significantly fewer parameters. The outperformance of TAPIS over TAPIS-VST in phases, steps, and actions recognition proves that our global video feature extractor is optimal and sufficient to understand the temporal progressions of surgical procedures. Although TAPIS-VST yields a better instrument region classification performance due to its finer-grained spatial processing and its increased amount of parameters, it still surpasses TAPIS by less than a 0.5% relative improvement with more than twice the parameters, hence highlighting TAPIS’s satisfactory performance.
@@
Moreover, Table 4 also compares TAPIS’ performance with the TAPIR model (Valderrama et al. (2022)). To ensure a fair comparison, we train TAPIR with its original implementation details, parameters, and box proposal architecture (Deformable DETR (Zhu et al. (2021))) on the training set of GraSP and test its performance on the test set. This comparison reveals that TAPIS significantly outperforms TAPIR across all tasks in the GraSP benchmark, especially in the novel and highly challenging atomic action detection task, where TAPIS achieves more than a 53% relative improvement over TAPIR’s results. The performance difference proves the potential of our instrument segmentation annotations for short-term tasks and underscores the impact of our technical improvements and methodological experimentation. Notably, TAPIR exhibits a slight decrease in performance for short-term tasks compared to the originally reported metrics in PSI-AVA [^11] https://github.com/BCV-Uniandes/TAPIR, which we attribute to the refining and curation of our dataset that eliminated biases and included annotations for challenging tiny instrument segments that were previously unannotated. The superiority of TAPIS over TAPIR and all alternative baselines positions it as state-of-the-art in all tasks of our GraSP dataset.
5.1 Validation of TAPIS in Public Benchmarks
In this section, we validate the consistency of our proposed methods by analyzing TAPIS’s performance on multiple publicly available benchmarks. Specifically, we compare the performance of our method with either previously proposed models on those benchmarks or our own proposed alternative baselines. In turn, these comparisons lend themselves to assessing our dataset’s solidity by contrasting the relative performance behavior of our proposed architectures in alternative benchmarks to their performance behavior in GraSP. Thus, Subsection 5.1.1 describes how we follow the conventional instrument and action recognition formulation in Twinanda et al. (2016); Nwoye et al. (2022); Maier-Hein et al. (2021); Wagner et al. (2023) and adapt our benchmark and methods for frame-wise recognition of instrument and action presence. Later, Subsections 5.1.2 and 5.1.3 explain how we also tailor our methods to analogous datasets in phases, steps, and gesture recognition (Huaulmé et al. (2021); van Amsterdam et al. (2022)), and even instrument-centric action localization (Islam et al. (2020)).
@@ 5.1 Validation of TAPIS in Public Benchmarks
Model Presence $mAP$ (%)
@@
Rendezvous (Nwoye et al. (2022)) 80.41 20.65
TAPIS w/o region proposals 61.39 12.27
TAPIS (Ours) 95.05 43.57
5.1.1 Instruments and Actions Presence Recognition
Given that GraSP includes instrument localization annotations, we expand our model evaluation to cover simpler and coarser tasks, such as recognizing the presence of instruments and associated actions. For this purpose, we adapt the dataset by transforming annotations into single binary vectors of dimensions $[1,N]$ for multi-label classification in each frame, where $N$ represents the number of classes (7 for instruments presence and 14 for actions presence). To enable TAPIS to execute this task, we modify the model output to consolidate predictions from each predicted region into a single classification vector. Since it is our best region proposal method, we employ our instrument segmentation baseline with a SwinL backbone as the region proposal model.
We compare TAPIS’s performance with two alternative baselines to validate our data and model. First, an alternative implementation of TAPIS without region proposal features that predicts instruments and actions presence directly from the video feature extractor. Second, with Rendezvous (Nwoye et al. (2022)), a state-of-the-art model for recognizing the presence of actions and instruments. This model predicts triplets in the form of $<$Instrument, Verb, Target$>$, where verb categories indicate the actions of each instrument in the keyframe, treated as analogous to the action categories in GraSP. To adapt Rendezvous for instrument presence recognition, we utilize all instrument categories, along with the “null verb” and “null target” categories, forming the triplets. Similarly, for the action presence task, we employ all action categories, along with the “instrument” and “null target” categories, to create the triplets. Additionally, we modify Rendezvous’ outputs according to the multi-label nature of GraSP actions. This adjustment included generating triplets for all actions present within each keyframe. Importantly, we train the models without pretraining on surgical datasets to ensure a fair comparison while preserving the pretraining in COCO of our region proposal network. The evaluation metric used for assessing performance was the $mAP$.
The results in Table 5 demonstrate that TAPIS strongly outperforms its implementation without region proposals and the Rendezvous model. These findings underscore the significance of having localized information from instruments and actions over time, as they provide valuable insights for the model and enhance its overall understanding. The results further validate that the spatial information offered by GraSP is essential for a comprehensive understanding of the surgical scene. These outcomes are also influenced by the temporal component of TAPIS, enabling more consistent predictions over time and integrating global information from the video feature extractor with local information from the region proposal method.
Table 6: Comparative results of TAPIS in Instrument and Interactions Detection on the EndoVis 2018 dataset. We test our TAPIS model in the benchmark proposed in Islam et al. (2020) and compare our performance with previously proposed methods. The best results are shown in bold.
@@ 5.1.2 Interactions Recognition in Endovis 2018
To compare our method with other models in alternative instrument and action detection benchmarks, we test our model on the Endovis 2018 interaction dataset (Islam et al. (2020)) and compare its performance with a current publicly available state-of-the-art method for this dataset. We adhere to the established splits in (Islam et al. (2020)), excluding the stapler class present in only one sequence of the dataset. Additionally, we omit the kidney class during the training and evaluation of TAPIS, as our model does not include tissue information to predict the actions performed by each instrument. In this experiment, we train TAPIS to simultaneously address instrument and action detection by combining the instrument and action losses multiplied by fixed coefficients. The composite loss function used for training used a 0.95 coefficient for the action detection loss and a 0.05 coefficient for the instrument detection loss.
We employ DINO (Zhang et al. (2023)) as our instrument detector and use the proposed $mAP@0.5IoU_{box}$ metric for evaluation. For a fair comparison, we evaluate the predictions made by the other models for each instrument (excluding the kidney category) as a conventional detection problem. We compare our method with MSLRGR-GISFSG (Seenivasan et al. (2022)), which simultaneously performs semantic instrument segmentation and interaction detection utilizing a graph reasoning approach. To compare accurately with this method, which relies on ground-truth bounding boxes for inference, we assume perfect bounding box locations for instruments, as this model does not perform direct detection. For interactions, we used the model’s predicted class for each box. We assign the bounding box class for instruments based on the instrument mask predicted within it.
On the one hand, the results shown in Table 6 highlight the significant proficiency of our method in instrument localization, once more emphasizing the superiority of transformer-based models for detection tasks compared to CNN approaches. On the other hand, TAPIS outperforms a state-of-the-art method in the interaction detection task. Despite the graph structure of the dataset, TAPIS effectively extracts rich information from each keyframe by combining global and temporal information from the video feature extractor and local details from the detector, proving that our proposed methodology is highly competitive in comprehending complex relationships between instruments with the surgical scene.
@@ 5.1.3 Gesture Recognition in RARP45
Table 7: Comparative results of TAPIS in gesture recognition task on the RARP-45 dataset. We compare the performance of TAPIS and its alternative baseline architectures on gesture recognition in the dataset proposed by van Amsterdam et al. (2022). The best results are shown in bold.
Method FLOPs (G) Param (M) Gesture Recognition $mAP$
SlowFast 25.29 33.7 49.00
TAPIS-VST 65.76 87.7 52.65
TAPIS 70.80 36.3 57.25
We further validate our methodology using the RARP-45 public gesture recognition benchmark. Given the absence of a public test set in the original RARP-45 dataset, we create our splits, using 27 videos for training and 9 videos for evaluation. To ensure the integrity of the splits, we maintain a consistent class distribution with the original training set. Specifically, for the test split, we select videos 2, 3, 4, 6, 7, 9, 14, 17, 23, 28, 32, and 33, while we use the remaining videos for the training split.
Following our previous experimental framework, we compare our method with its alternative versions using SlowFast and VST as video feature extractors. All experiments employed a window of 16 frames with a stride of 1. As shown in Table 7, our TAPIS model outperforms the alternative baselines, with all transformer-based methods surpassing the CNN model. This behavior demonstrates our model’s robustness by achieving outstanding performance in tasks analogous to the ones proposed in GraSP yet differently formulated. These results are consistent with the findings in the GraSP dataset, thus reinforcing the effectiveness of our approach in handling video-level tasks and proving the reliability of our results.
@@ 5.1.4 Phase and Step Recognition in MISAW
To further validate our results in phases and steps recognition, we test our models on the MISAW dataset (Huaulmé et al. (2021)), which is analogous in long-term tasks to GraSP. We again follow the experimental setup we define for GraSP and compare the performance of the different explored architectures for GraSP (SlowFast and TAPIS-VST) in the phases and steps recognition benchmarks of the MISAW dataset.
We use a temporal window of 16 frames and a temporal stride of 1 second for our setting in MISAW. Moreover, we use cases 1-17 for training and 18-27 for validation, as proposed by the authors in Huaulmé et al. (2021). The results in Table 8 portray the performance of our TAPIS model when evaluated on the MISAW dataset. Compared to both SlowFast and VST architectures, TAPIS performs better on both phases and steps recognition tasks. These results are also consistent in proving the potential advantages of using Transformers for video feature extraction over CNNs across both benchmarks and rectifying the correctness of the method we propose for addressing phase and step understanding tasks.
Table 8: Comparative results of TAPIS in phases and steps recognition on the MISAW dataset. We compare the performance of TAPIS and its alternative baseline architectures on Phase and Step recognition in the benchmark proposed by Huaulmé et al. (2021). The best results are shown in bold.
Method FLOPs (G) Param (M) Long-Term Tasks $mAP$ (%)
Phases Steps
SlowFast 25.29 33.66 94.04 70.50
TAPIS-VST 65.76 87.65 96.16 73.46
TAPIS 70.80 36.31 97.14 77.52
@@ 5.2.1 Long-Term Tasks Experiments
We conduct design studies to determine the optimal temporal window for both phases and steps recognition tasks. We aim to define the best configuration for each task within our computational resource limitations. We present our results in Tables 9 and 10.
As stated earlier in this work, phases comprise broader temporal segments within surgical procedures, implying the need for a broader temporal context to recognize them effectively. In our phases recognition temporal window design experiments illustrated in Table 9, we validate that a wide-spaced temporal window correlates positively with the phases recognition performance. These findings substantiate the statement that temporal windows that allow wide-reaching temporal context capture more relevant information for accurate phases recognition. However, the refinement of this temporal context showcases a trade-off in model performance, where excessively large and spaced temporal windows tend to hinder performance. Consequently, our optimal temporal context for accurate phase recognition is a 16-frame window with a 4-second temporal stride, capturing 64 seconds centered on the frame of interest. Remarkably, our optimal window design is consistently related to average phase duration statistics showcased in Figure 4, where phase duration ranges mostly from 45 to 300 seconds approximately. Therefore, a 64-second temporal window encapsulates the semantic context needed to discern among different surgical phases.
In contrast, steps represent finer-grained time segments within these phases. Consequently, discerning these discrete steps requires a more detailed temporal context around the central frame to extract relevant features for step differentiation effectively. As detailed in Table 10, our analysis highlights that employing temporal windows sampled with shorter temporal stride (higher frame rates) results in higher performance in steps recognition. Increasing the temporal stride reduces performance, and excessively augmenting it yields the lowest $mAP$ values observed. We find that the optimal temporal window for step recognition is to keep a 16-frame window size and use a low 1-second temporal stride so the model captures a temporal window of 16 seconds centered on the frame of interest. Our optimal window design is consistently related to average step duration statistics from Figure 5, where step duration ranges mostly from 10 to 200 seconds approximately. Notably, steps exhibit a wider range of durations compared to phases. This variability implies that employing a uniform temporal window to identify all step categories poses a challenge, as certain instances might require a larger temporal context for accurate identification compared to others.
Studying the temporal windows for each task allows us to acknowledge the relationship between the temporal window resolution designed for each task and the multi-level understanding nature of our dataset. As we show in Table 9 and 10, tasks with finer granularity, such as steps, require higher temporal resolutions for our model to discern individual steps effectively. In contrast, tasks with lower granularity, such as phases, demand lower resolutions of temporal context for our model to distinguish them accurately.
Table 9: Temporal window design experiments for Phase Recognition in the cross-validation set. The number of frames denotes the temporal window size centered on the frame of interest. Temporal stride is the gap between consecutive sampled frames of the sequence in seconds. The best results are shown in bold.
Window Size Temporal Stride (seconds) $mAP$ (%)
16 1 72.21 ±$\pm 0.8$
16 4 72.90 ±$\pm 1.65$
16 8 71.40 ±$\pm 0.24$
24 1 72.41 ±$\pm 1.04$
24 4 72.27 ±$\pm 1.22$
32 4 72.04 ±$\pm 1.70$
Additionally, we present more specific results of phases and steps recognition per-class performance in the Supplementary Material. Initially, the assessment of the model’s performance across diverse phase categories reveals differing levels of success in identifying individual surgical phases. Notably, phases 1-6 and 10 consistently attain relatively higher $AP$ scores across all three methods, while categories 7 and 8 exhibit relatively lower scores. When correlating these observations with the dataset statistics (Figure 2), it is notable that the phases with lower $AP$ scores across the methods display a low representation in the dataset, as it is natural when studying long-tailed datasets. Additionally, the lower mean duration of these phases (Figure 4) might contribute to the model finding it more challenging to discern and accurately predict these phases, such as classes 7 and 8. This relation is related to the broader temporal stride that may not be optimal for shorter phase instances that span fewer consecutive frames.
Table 10: Temporal window design experiments for Step recognition in the cross-validation set. The number of frames denotes the temporal window size centered on the frame of interest. Temporal stride is the gap between consecutive sampled frames in seconds. The best results are shown in bold.
Window Size Temporal Stride (seconds) $mAP$ (%)
16 0.10 45.40 ±$\pm 0.78$
16 0.23 46.56 ±$\pm 0.66$
16 1 48.36 ±$\pm 1.14$
16 4 43.95 ±$\pm 1.07$
24 0.10 45.31 ±$\pm 0.17$
24 0.23 46.55 ±$\pm 0.07$
24 1 48.23 ±$\pm 2.30$
32 0.10 45.13 ±$\pm 0.36$
32 0.23 46.55 ±$\pm 0.44$
32 1 48.02 ±$\pm 1.13$
@@
$mAP@0.5IoU_{box}$ $mAP@0.5IoU_{segm}$ mIoU IoU mcIoU
TernausNet (Shvets et al. (2018)) UNet 16 - - 41.74 ±$\pm 5.07$ 24.46 ±$\pm 6.04$ 16.87 ±$\pm 3.70$
MF-TAPNet  (Jin et al. (2019)) TAPNet - - 66.63 ±$\pm 1.24$ 29.23 ±$\pm 1.43$ 24.98 ±$\pm 0.59$
ISINet  (González et al. (2020)) R50 201 44 79.85 ±$\pm 2.17$ 78.29 ±$\pm 2.82$ 78.44 ±$\pm 1.13$ 70.85 ±$\pm 0.00$ 56.67 ±$\pm 1.46$
QPD Mask DINO (Dhanakshirur et al. (2023)) R50 228 54 88.46 ±$\pm 1.14$ 87.39 ±$\pm 1.75$ 83.89 ±$\pm 1.23$ 82.56 ±$\pm 1.14$ 74.36 ±$\pm 1.04$
Similarly, the model’s performance across various step categories reveals that steps 12 to 18 and 20, characterized by lower mean durations, exhibit lower Average Precision ($AP$) scores across the evaluated methods. Additionally, less represented step classes in the dataset, like 6, 7, and 11, tend to manifest lower $AP$ scores, and steps 0, 1, and 19, with higher representation, demonstrate comparatively better performance. However, the model’s performance was still low for some highly represented classes like 2, 3 and 5. This effect may be due to the high variability in these steps’ duration, as shown in Figure 5. Once again, the varying performance between classes is related to the fixed temporal window design, which may not be optimally suited to capture the high inter and intra-class variability in the duration of different step instances.

### v3

@@ 5 Experiments and Results
Following the experimental setup in [84], we compare TAPIS with two alternative models: SlowFast ([25]) as a CNN-based baseline using ISINet ([32]) for region proposals, and an alternative implementation called TAPIS-VST, which replaces TAPIS’s video feature extractor with the Video Swin Transformer (VST) ([56]). The results in Table 4 demonstrate that both TAPIS and TAPIS-VST outperform the CNN-based baseline. This finding highlights the advantages of fully transformer-based approaches, as their enhanced modeling capabilities, powered by long-range attention mechanisms, improve the understanding of complex surgical scenes. Additionally, TAPIS surpasses TAPIS-VST in most tasks while achieving comparable instrument segmentation performance with significantly fewer parameters. We attribute TAPIS-VST’s slight advantage in instrument region classification to its finer-grained spatial processing and a larger number of parameters.
@@
Table 4 compares TAPIS with the TAPIR model ([84]), which we trained using its original implementation details and box proposal network on GraSP’s training set. TAPIS significantly outperforms TAPIR across all tasks, achieving a 53% relative improvement in the challenging atomic action detection task. This performance difference validates the impact of our enhanced instrument segmentation annotations and our technical and methodological advancements. Notably, TAPIR exhibits a slight decrease in performance for short-term tasks compared to the initially reported metrics [^8] https://github.com/BCV-Uniandes/TAPIR, which we attribute to the curation of our dataset that eliminated biases and included annotations for previously unannotated and challenging tiny instrument segments.
In the following sections, we validate TAPIS by comparing it with publicly available state-of-the-art (SOTA) models designed for specific single tasks similar to those in GraSP. For this purpose, we evaluate TAPIS, its alternative baselines, and these SOTA models on our cross-validation set and on alternative benchmarks featuring tasks analogous to GraSP. To ensure a fair comparison, we train each SOTA model independently for its respective task and adapt all temporal models for offline processing to align with TAPIS’s framework. These comparisons also assess the consistency of our GraSP dataset by analyzing how the performance ranking of various models on external benchmarks aligns with their performance on GraSP and the expected performance ranking from the current SOTA.
@@
Model Regions Time Presence $mAP$ (%)
@@
Rendezvous ([65]) 80.41 20.65
RiT ([76]) 78.20 21.76
TAPIS (Ours) 53.19 24.77
84.76 41.26
93.54 31.04
94.33 43.85
Table 6: Comparative results of TAPIS in Instrument and Interactions Detection on the EndoVis 2018 dataset. We test our TAPIS model in the benchmark proposed in [39] and compare our performance with previously proposed methods. The best results are shown in bold.
@@
5.1 Validation of TAPIS in Public Benchmarks
5.1.1 Instruments and Actions Presence Recognition
We adapt our benchmark and methods for simpler, frame-level tasks such as instrument and action binary presence recognition, following the conventional formulations in [83, 65, 60, 87]. We modify TAPIS for this task by using our best instrument segmentation baseline and consolidating predictions from all proposed regions into a multi-label binary classification vector.
We compare TAPIS with two groups of baseline methods for this task. First, we adapt Rendezvous ([65]) and RiT ([76]) on our benchmark by combining all instrument categories, along with the ”null verb” and ”null target” placeholders and using all action categories combined with the ”instrument” and ”null target” placeholders. Consequently, we train Rendezvous and RiT to recognize the presence of instruments and actions in our keyframes. Second, we test alternative implementations of TAPIS without region proposal features, by directly predicting instrument and action presence from the video feature extractor; and without temporal processing by running TAPIS on a single frame. We train all models without pretraining on any surgical datasets to ensure a fair comparison, and we assess performance using the $mAP$.
The results in Table 5 demonstrate that TAPIS significantly outperforms its implementations without region proposals or temporal processing and achieves superior performance compared to the Rendezvous and RiT models for both instrument and action presence recognition. These results underscore the value of GraSP’s spatial annotations, as including localized information about instruments and actions enhances the model’s ability to understand and process the surgical scene. Additionally, TAPIS’s temporal component further improves recognition by incorporating valuable temporal context that enables the model to better capture dynamic interactions and temporal progressions within surgical workflows.
Table 7: Comparative results of TAPIS on the gesture recognition task of the RARP-45 dataset. We compare the performance of TAPIS with SOTA models for phase recognition, and its alternative baseline architectures on gesture recognition in the dataset proposed by [4]. The best results are shown in bold.
Method FLOPs (G) Param (M) Gesture Recognition (%)
mAP F1 score
TeCNO 4.50 23.64 43.08 45.94
Trans-SVNet 11.96 23.67 42.15 42.85
SAHC 11.95 25.73 43.79 41.80
SlowFast 25.29 33.67 49.00 46.79
TAPIS-VST 65.76 87.66 52.65 48.25
TAPIS (Ours) 71.80 36.31 57.25 54.17
Table 8: Comparative results of TAPIS in phases and steps recognition on the MISAW dataset. We compare the performance of TAPIS with SOTA models for phase recognition, and its alternative baseline architectures on Phase and Step recognition in the benchmark proposed by [36]. The best results are shown in bold.
Method FLOPS (G) Params. (M) Phases (%) Steps (%)
mAP F1 score mAP F1 score
TeCNO 4.50 23.64 95.58 62.83 68.89 62.99
Tras-SVNet 11.96 23.67 90.38 88.17 72.44 63.07
SAHC 11.95 25.73 95.10 92.43 78.79 67.88
SlowFast 25.29 33.67 94.04 81.98 70.50 62.48
TAPIS-VST 65.76 87.66 96.16 89.71 73.46 63.36
TAPIS 70.80 36.31 97.14 90.12 77.52 66.51
Table 9: Comparative results of TAPIS in phase and step recognition on GraSP’s cross-validation set. We evaluate the TAPIS trained for phase and step recognition jointly (multi-task) and independently, and compare it with alternative baseline and SOTA models for surgical phase recognition trained independently. The best results are shown in bold.
Model Multi-Task FLOPs (G) Params. (M) Phase Recognition (%) Step Recognition (%) Long-Term Average (%)
$mAP$ F1 score $mAP$ F1 score $mAP$ F1 score
TeCNO ([15]) 4.50 23.64 64.64 ±$\pm 2.88$ 61.51 ±$\pm 2.73$ 27.39 ±$\pm 1.78$ 22.70 ±$\pm 2.21$ 46.01 ±$\pm 21.59$ 42.11 ±$\pm 22.50$
Tran-SVNet ([28]) 11.96 23.67 66.89 ±$\pm 1.79$ 58.20 ±$\pm 3.80$ 33.96 ±$\pm 2.02$ 30.31 ±$\pm 2.22$ 50.42 ±$\pm 19.08$ 44.25 ±$\pm 16.30$
SAHC ([19]) 11.95 25.73 70.80 ±$\pm 1.76$ 66.35 ±$\pm 1.60$ 37.55 ±$\pm 1.08$ 35.85 ±$\pm 1.24$ 54.17 ±$\pm 19.23$ 50.10 ±$\pm 18.80$
SlowFast 25.29 33.67 68.44 ±$\pm 0.10$ 58.59 ±$\pm 3.06$ 43.72 ±$\pm 1.79$ 40.29 ±$\pm 3.81$ 56.15 ±$\pm 14.23$ 49.44 ±$\pm 10.94$
TAPIS-VST 65.76 87.66 70.46 ±$\pm 0.09$ 62.25 ±$\pm 0.33$ 45.51 ±$\pm 1.48$ 39.45 ±$\pm 0.87$ 57.98 ±$\pm 14.43$ 50.85 ±$\pm 13.17$
65.76 87.68 70.60 ±$\pm 1.59$ 61.64 ±$\pm 3.32$ 47.17 ±$\pm 2.65$ 42.25 ±$\pm 3.72$ 58.89 ±$\pm 13.64$ 51.95 ±$\pm 11.56$
TAPIS 70.80 36.31 72.31 ±$\pm 2.04$ 64.04 ±$\pm 2.13$ 49.72 ±$\pm 2.35$ 45.57 ±$\pm 2.03$ 61.01 ±$\pm 13.17$ 54.81 ±$\pm 10.80$
70.80 36.32 71.36 ±$\pm 1.33$ 64.72 ±$\pm 1.92$ 50.74 ±$\pm 2.53$ 46.91 ±$\pm 3.36$ 61.05 ±$\pm 12.02$ 55.82 ±$\pm 10.52$
@@ 5.1.2 Interactions Recognition in Endovis 2018
We evaluate TAPIS on the EndoVis 2018 interaction dataset ([39]), following its established splits. We exclude the stapler class, which only appears in a single sequence, and omit the kidney class, as TAPIS does not incorporate tissue information. For this experiment, we replace Mask2Former with DINO ([91]) as the region proposal network, trained for instrument detection using the annotations from [39]. Additionally, we train TAPIS without pretraining the video feature extractor on surgical datasets. TAPIS is then jointly trained for instrument and interaction detection by combining their respective losses, weighted at 0.95 for action loss and 0.05 for instrument loss.
We compare TAPIS with MSLRGR-GISFSG ([74]) by evaluating their instrument and action predictions using the $mAP@0.5IoU_{box}$ metric. Since MSLRGR-GISFSG relies on ground-truth bounding boxes for inference, we assume perfect bounding box locations to evaluate its performance, assigning instrument classes based on predicted masks and assessing interactions using each bounding box’s predicted interaction with the kidney.
As shown in Table 6, TAPIS significantly outperforms MSLRGR-GISFSG in both instrument localization and interaction detection tasks, proving the robustness of our transformer-based approach for surgical localization tasks. Despite the dataset’s graph structure, TAPIS effectively integrates global features from the video feature extractor with local detections from DINO, demonstrating its capability to model complex instrument-scene relationships across diverse datasets.
@@ 5.1.3 Gesture Recognition in RARP45
We validate TAPIS on the RARP-45 benchmark for gesture recognition. Given the absence of a public test set in the original RARP-45 dataset, we create data splits as detailed in D of the Supplementary Material. Following our proposed experimental framework, we train our video feature extractor on RARP45 and compare it with SlowFast, TAPIS-VST, and multiple SOTA models designed for surgical phase recognition ([15, 28, 19]). Table 7 shows that TAPIS outperforms alternative baselines, with all transformer-based methods achieving the best results. Similarly, our model strongly surpasses SOTA models for phase recognition due to the finer-grained nature of the gesture recognition task for which these architectures have not been designed. This behavior shows TAPIS’s adaptability to tasks analogous to those proposed in GraSP yet differently formulated.
@@ 5.1.4 Phase and Step Recognition in MISAW
To further validate our phase and step recognition benchmark, we evaluate TAPIS on the MISAW dataset ([36]) using MISAW’s predefined dataset splits and follow the comparative experimental setup described in Section 5.1.3. The results in Table 8 show that TAPIS outperforms most SOTA models and its alternative baselines, reinforcing the advantages of transformers for surgical video analysis and demonstrating the consistency of our approach.
Nevertheless, TAPIS slightly underperforms SAHC, and we attribute this to MISAW’s simpler formulation of phase and step recognition tasks, which are characterized by fewer categories, shorter simulated videos, and longer, uninterrupted temporal segments. In contrast, GraSP includes multiple phase and step categories with highly variable and interrupted segment durations on long in-vivo videos. Thus, models like SAHC, which process extended temporal contexts of global frame-wise features and are tailored for datasets with similar phase recognition setups (e.g., Cholec80 ([83]), M2CAI16 ([79])), may be better suited to MISAW’s data. Nonetheless, despite being optimized for GraSP’s more complex setup, TAPIS achieves competitive performance and demonstrates its generalization ability across diverse benchmarks.
@@ 5.2.1 Long-Term Tasks Experiments
Table 9 presents the performance of TAPIS and TAPIS-VST trained for phase and step recognition, independently and jointly. To directly compare models, we calculate the average Long-Term performance as the average per-metric across all tasks and folds. Even when trained independently, TAPIS outperforms alternative models on both tasks and demonstrates better long-term performance, while only SAHC achieves comparable performance in phase recognition. Additionally, TAPIS shows a significant advantage over SOTA models in step recognition, thus highlighting limitations in phase recognition models, which lack versatility for more granular tasks. In contrast, TAPIS’s single-stage, video-wise encoder captures finer spatio-temporal details that allow better performance in finer temporal tasks. Similarly, the relative performance ranking of the SOTA models aligns with current phase recognition benchmarks, confirming our benchmark’s validity. Finally, TAPIS yields better results when trained jointly for phase and step recognition, with the most notable improvements observed in the challenging step recognition task. This demonstrates TAPIS’s multi-task learning capabilities, allowing it to leverage task correlations and perform both tasks within a single training run.
$mAP@0.5IoU_{box}$ $mAP@0.5IoU_{segm}$ mIoU IoU mcIoU
TernausNet ([78]) UNet 16 - - 41.74 ±$\pm 5.07$ 24.46 ±$\pm 6.04$ 16.87 ±$\pm 3.70$
MF-TAPNet  ([41]) TAPNet - - 66.63 ±$\pm 1.24$ 29.23 ±$\pm 1.43$ 24.98 ±$\pm 0.59$
ISINet  ([32]) R50 201 44 79.85 ±$\pm 2.17$ 78.29 ±$\pm 2.82$ 78.44 ±$\pm 1.13$ 70.85 ±$\pm 0.00$ 56.67 ±$\pm 1.46$
QPD Mask DINO ([18]) R50 228 54 88.46 ±$\pm 1.14$ 87.39 ±$\pm 1.75$ 83.89 ±$\pm 1.23$ 82.56 ±$\pm 1.14$ 74.36 ±$\pm 1.04$
@@
We present per-class performance results for phase and step recognition in Tables 10 and 11 of the Supplementary Material. For phases, classes 1–6 and 10 consistently achieve higher AP scores, while classes 7 and 8, characterized by lower representation in the dataset and shorter mean durations, exhibit lower scores. Similarly, step categories with lower durations, such as steps 12–18, or those with sparse representation, like steps 6, 7, and 11, show lower AP scores, while steps 0, 1, and 19 perform well due to higher representation. Notwithstanding, some highly represented classes like steps 2, 3, and 5 still underperform, likely due to high variability in their durations.